Please help me, Please.
A retired couple has $130,000 to invest to obtain annual income. They want some of it invested in safe Certificates of Deposit yielding 5%. The rest they want to invest in AA bonds yielding 10% per year. How much should they invest in each to realize exactly %10,500 per year?

Answer by stanbon(75887)

A retired couple has $130,000 to invest to obtain annual income. They want some of it invested in safe Certificates of Deposit yielding 5%. The rest they want to invest in AA bonds yielding 10% per year. How much should they invest in each to realize exactly %10,500 per year?
-----------------
Let amt. invested at 5% be "x" ; interest on this is 0.05x dollars
------------
Amt. invested at 10% is "130,000-x"; interest is 0.10(130000-x)=13000-0.1x dollars
-----------------
EQUATION:
interest + interest = 10,500
0.05x + 13000-0.10x = 10500
-0.05x = -2500
x = $50,000 (Amount they should invest in CD's at 5%)
130000-x = $80,000 (Amount they should invest in AA bonds at 10%)
===============
Cheers,
Stan H.